OpenStudy (anonymous):

The function below shows the number of students of a school who enrolled for ballet classes. Let f(x) represent the total number of students who enrolled for the classes after x years. f(x) = 12(1.06)^x The average rate of change in the number of students who enrolled for ballet classes from the third to the fifth year is ________students per year. Round your answer to the nearest whole number.

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

I think the answer is 1.

OpenStudy (solomonzelman):

First find the change. 12(1.06)^5 - 12(1.06)^3 and divide by 2 I think this is it.

OpenStudy (anonymous):

f(x) = 12(1.06)^x f(x) = 12(1.06)^3 f(x) = 12(1.191016) f(x) = 14.292192 f(x) = 12(1.06)^x f(x) = 12(1.06)^5 f(x) = 12(1.33822558) f(x) = 16.058707 Average Rate of Change: m = y2 - y1/x2 - x1 Points: (3,14.292192) and (5,16.058707) m = 16.058707 - 14.292192/5 - 3 m = 1.766515/ 2 m = 0.8832575 So the average rate of change in the number of students who enrolled for ballet classes from the third to the fifth year is approximately 1 student per year.

OpenStudy (anonymous):

12(1.06)^5 - 12(1.06)^3/ 2 = 0.8832575 So I got the same thing you did.

OpenStudy (anonymous):

Thanks :)

OpenStudy (solomonzelman):

Anytime!

OpenStudy (solomonzelman):

I just thought of it as average of numbers, (HAte thinking about functions)

OpenStudy (anonymous):

Lol okay

OpenStudy (solomonzelman):

Cool, so we are clear on this, right?

OpenStudy (anonymous):

Yeah

OpenStudy (solomonzelman):

Good!

OpenStudy (anonymous):

I got the same thing but wasn't sure if it was right because it asked for you to round to the nearest whole #